A rectangular fish tank has a base 2 feet wide and 3 feet long. When the tank is partially filled with water, a solid cube with an edge length of 1 foot is placed in the tank. If no overflow of water from the tank is assumed, by how many inches with the level of the water in the tank rise when the cube becomes completely submerged?

Answer by mananth(16949)

A rectangular fish tank has a base 2 feet wide and 3 feet long. When the tank is partially filled with water, a solid cube with an edge length of 1 foot is placed in the tank. If no overflow of water from the tank is assumed, by how many inches with the level of the water in the tank rise when the cube becomes completely submerged?
The base of the tank is 3ft by 2 ft.
If there is 1 foot water in the tank the total volume = 3*2 * 1
= 6 cubic feet.
volume of 6 cu. feet coressponds to 1 foot of water in height.
..
The volume of the immersed solid cube is 1*1*1 = 1 cu.ft
.
The volume increase in the water is equal to the volume of immersed solid.
for every 6cu.ft of water the height is 1 foot or 12 inches
for 1 cu.ft of water the height will be 12/6 = 2 inches
The height will increase by 2 inches.